solve log 2x+log 10=3
Answer by lwsshak3(11628)

solve log 2x+log 10=3
..
log 2x+log 10=3
log(2x*10)=3 (log multiplication rule)
change to exponential form:(base(10) raised to log of the number(3)=number(2x*10)
10^3=2x*10
1000=20x
x=50
check: log(2*50)+log (10)=log 100+log 10=3